# Bayesian Quantum Amplitude Estimation

### v1

@@ II.6 Literature overview
Key idea Parallelizable Circuits Main strength(s) Main weakness(es) Complexity
QAE
[1]
phase estimation on Grover operator no QAE + QPE provably
optimal complexity
circuit depth and width; noise oblivious
$N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha}\cdot a^{-1})$
MLAE
[18]
heuristic circuits $\rightarrow$ statistical estimation fully QAA simplicity; solid numerical performance no formal guarantees; depth grows indefinitely best case $\mid$ observed: $N_{q}^{\text{(LIS)}}\sim\epsilon^{-0.75}\mid\epsilon^{-0.76}$
$N_{q}^{\text{(EIS)}}\sim\epsilon^{-1}\mid\epsilon^{-0.88}$
V-MLAE
[35]
MLAE with variational approximation partially QAA numerical performance; limits circuit depth no formal guarantees; cost overhead similar to MLAE except for EIS-observed (untested)
AES
[16]
rough estimate $\rightarrow$ exponential refining partially QAA provably
optimal complexity
large quantum cost offset; noise oblivious $N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha}\cdot a^{-1}$ )
SAE
[17]
amplification
$\rightarrow$ probability inversion
partially Hadamard tests simple non-essential circuits; inconclusive analysis $N_{q}\in\mathcal{O}(a^{-1})$
IAE
[21]
watchful optimization of Fisher inf. partially QAA provably near optimal complexity depth grows indefinitely; noise oblivious
M-IAE
[27]
IAE but distribute shots more favorably partially QAA provably
optimal complexity
same as IAE
$N_{q}\in\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\alpha})$
FAE
[20]
complementary measurements $\rightarrow$ invert probability partially QAA provably near optimal complexity depth grows indefinitely; noise oblivious
BAE
(this paper)
problem tailored Bayesian inference partially QAA optimal observed complexity, noise resilient limited formal guarantees; classical cost $N_{q}\in\mathcal{O}(\epsilon^{-1})$ (observed)
III Bayesian amplitude estimation
@@
Figure 2 summarizes the inference process applied to QAE. We use vector notation to emphasize sequences.
Figure 2: Diagrammatic representation of a Bayesian algorithm for quantum amplitude estimation.
@@
One alternative is to find the mean, which requires integrating over the distribution. We will discuss integration methods in subsection III.2. But before, we note that this extra step arises due to one of Bayesian inference’s strongest features. The mean is just a particular case of an expected value we can compute based on data. More generally, we can calculate the expectation for any function of the parameter(s),
@@
In figure 2, it would take the form of an a priori optimization of $\vec{m}$, the sequence of controls, considering expectations over the prior. However, the quality of the predictions is heavily dependent on the accuracy of the distribution; that is, how close it is to reality (a Dirac delta function on the real value). Thus, this choice may be unsatisfactory, especially in the case of uninformative priors.
Instead, we can unfold the inference process into several steps, considering intermediary distributions $\mathbf{P}(\theta\mid D_{0})$, $\mathbf{P}(\theta\mid D_{0},D_{1})$, and so on. The freedom to do so stems from the structure of Bayes’ rule: one may do a bulk update based on a complete dataset, or consider the data incrementally. In the second case, the posterior of one iteration becomes the prior for the next, allowing empirical knowledge to inform the latter’s controls. The inference process can then occur online and adaptively, exploiting all available information at each point. There are additional benefits to this approach, to be discussed in section III.3.
We adopt this adaptive approach for QAE. Figure 3 depicts the structure of such a scheme.
Figure 3: Diagrammatic representation of an adaptive Bayesian algorithm for quantum amplitude estimation.
Here we consider a single datum per iteration, but the process can be done with arbitrary granularity, dividing the measurements into blocks of several and considering a chunk of data per step. There are multiple trade-offs involved. Such details on the experimental design are discussed in section III.3 and appendix A, where we also introduce problem-tailored tricks to cut back on the optimization cost (the most resource-intensive sub-routine).
@@
This finalizes the main outline of the algorithm. Pseudo-code is provided in algorithm 1.
Algorithm 1 Algorithm for adaptive inference.
1: Parameters: missingwNshots
2:prior $\triangleright$ Distribution over parameter.
3:${\hat{\theta}}$triangleright Parameter estimate.
4:$m\leftarrow$ 0
[…]
7:while not termination_criterion do
[…]
11:end while
12:$\hat{\theta}\leftarrow$ estimator(d)
13:return ${\hat{\theta}}$
For clarity, we distinguish the algorithm hyperparameters from the input. The parameters can be tuned for the general case, while the input variables are problem dependent. The prior distribution is also problem dependent in that previous information may be available. However, for all numerical simulations performed herein, we assume the absence of prior knowledge, represented by a flat initial distribution over $a\in[0,1]$.
The termination criterion may be based on a number of iterations, a number of queries/probing time, or a target uncertainty.
The estimator is a construction mapping the posterior to an estimate - such as the mean, the mode or the median. The details of how the probability distribution is represented are left unspecified, and left for subsection III.2. This will affect the functions used for the Bayesian updates, and for the expected value calculations.
III.1 Noise tailoring
Another advantage of the Bayesian approach is its natural ability to accommodate noise models by adapting the likelihood function. This brings the ability to study the undesirable properties alongside the parameters of interest. As the former undermine the characterization of the latter, this offers a largely customizable path to robust characterization, without any structural modifications to the algorithm. On the one hand, we can account for the fact that the system’s behavior is affected by confounding factors, and correct the estimates accordingly. On the other, we are in a position to pursue realistic measurement selection under less than ideal conditions.
To enjoy the flexibility, robustness and scalability of integration, it is vital that one employs efficient integration methods. Furthermore, to enable adaptivity, these methods must be compatible with sequential data processing. Section III.2 discusses this.
In the simulations performed for this paper, we consider an exponential damping model for decoherence (subsection IV.2), and a constant coherence time. BAE learns this time in a pre-processing phase as described above. The process is similar to the approach used to learn the amplitude. Details on the estimation of noise in BAE are given in subsection III.3.
III.2 Integration methods
The previous section discussed the strengths of Bayesian inference, and how they are tied to the probabilistic depiction of information. It remains to discuss the specifics of this depiction, which are determinant for the realization of such strengths.

### v5

@@ 2.6 Literature overview
3 Bayesian amplitude estimation
(a) Non-adaptive Bayesian algorithm for quantum amplitude estimation.
(b) Bayesian quantum amplitude estimation (BAE). Pseudo-code is provided in Algorithm 1; blocks choose m and Bayesian update refer to algorithms 8 and 2, respectively.
Figure 2: Diagrams of Bayesian algorithms for amplitude estimation.
We propose Bayesian amplitude estimation (BAE), a Bayesian inference based algorithm following the hybrid scheme of Figure 1. In each iteration, a number of Grover steps $m$ is chosen classically, after which the corresponding quantum circuit is executed. The measurement outcome is fed into the classical processing unit, and this cycle repeats as necessary to achieve the intended uncertainty.
The Bayesian framework allows us to refine our (possibly null) initial knowledge about the amplitude, representing it by probability distributions. This description allows us to evaluate the expected utility (variance) of any $m$ in a look-ahead fashion, and thus optimize $m$ or sequences $\vec{m}$. We opt for a greedy algorithm - unitary look-ahead, where a single $m$ is optimized at a time -, to minimize optimization costs. For related reasons, we perform a classical warm-up phase, in the initial stage where information is scarce.
The workflow is represented diagrammatically in Figure 2(b), while pseudo-code is provided in algorithm 1. Overviews of Bayesian inference and Bayesian experimental design are provided in Appendices A and C respectively. The Bayesian updates and estimator calculations are the backbone our inference-based protocol; for efficient numerical representations of the probability distributions, we use sequential Monte Carlo (SMC) with Markov Chain Monte Carlo transitions, detailed in appendix B.
The crux of the algorithm is the choose m sub-routine. A greedy strategy would involve discrete optimization over an infinite domain, as $m$ is unbounded above. To remedy this, we develop a heuristic strategy to define the search range adaptively, based on window expansions (Appendix C.1). We start with a low upper bound for the search range; each time one of the highest possible controls allowed by the window is chosen in the optimization, we register a hit. When the number of hits achieves a pre-determined threshold, a window expansion is triggered: the former upper bound becomes the new lower bound, while the upper bound doubles. We then reset the hit counter to zero, and start again. This relies on the fact that Heisenberg-limited QAE algorithms have amplifications schedules where $m$ grows exponentially with the iterations [1, 15, 50]. While the window width still grows exponentially, we achieve constant cost by having the evaluations within be exponentially sparse.
Importantly, this Bayesian framework can easily be adapted to a noisy scenario, by adapting the generative model. We extend the Bayesian framework to learn noise parameters beyond the amplitude, obtaining an algorithm that can characterize device noise and adapt in real time. Details about the noise tailoring are provided in Appendix D.
Our algorithm is capable of exploiting prior knowledge, e.g if the amplitude is known to be small (Appendix A); can evaluate noise models for negligible extra cost (Appendix B.4); is well-suited for parallel execution, with largely independent tasks that can be distributed across both quantum and classical devices (Appendix I); amd features adjustable parameters that can be adapted based on practical trade-offs (Appendix J).
Algorithm 1 Algorithm for adaptive inference. Non trivial functions are defined in Appendix B.1.
1: Parameters: missingwNshots, Nshots
2:prior $\triangleright$ Distribution over parameter.
[…]
7:while not termination_criterion do $\triangleright$ E.g. target uncertainty.
8:  $\texttt{m}\leftarrow$ optimize_control(d)
9:  $\vec{m}\leftarrow$ repeat(m, Nshots)
10:  $\vec{o}\leftarrow$ measure($\vec{m}$)
11:  $\texttt{d}\leftarrow$ batch_bayesian_update(d, $\vec{m}$, $\vec{o}$)
12:end while
[…]
14:return $^{\hat{\theta}}$
3.1 Annealed Bayesian amplitude estimation
BAE is an inference algorithm that greedily minimizes the variance. In this section, we propose a different approach - annealed BAE (aBAE), borrowing from annealed importance sampling [37, 47], a well-known algorithm in the statistics literature.
The successive powers of the distribution form a sequence of distributions that can be sampled from using sequential Monte Carlo. Note that previously, the sequence of distributions was given by cumulative datasets of increasing size. The SMC algorithm can be applied to any sequence of distributions. Refer to Appendix B.1 for details.
The advantage of this method is statistical robustness, which we expect to be especially notable in higher dimensions, where adequately exploring the space is more challenging. Additionally, in multi-dimensional estimation,the variance cost function must be generalized to a scalar representation of the covariance. This raises new points to consider, such as the relative scales of the parameters. Using the ESS eliminates these problems. Lastly, the complexity of calculating covariance-related metrics scales with the dimension, whereas the complexity of calculating the ESS does not.
4 Methods
4.1 Processing and benchmarking
For a graphical depiction of this assessment, we will represent the root mean squared error (RMSE) as a function of the number of queries $N_{q}$. In previous QAE proposals, tests are performed for a fixed amplitude [21, 36, 56] chosen by the authors. Such choices may lead to biased assessments, as well as overfitting.
, where $a_{i}$ and $\hat{a}_{i}$ are the real and estimated amplitude values, and $N$ is the number of random amplitudes used.
Similarly, we use a normalized version of the average standard deviation. This allows for a more thorough performance assessment, reveals behavior irregularities in some algorithms, and allows for universal tuning of the algorithm hyperparameters.
Importantly, we use a different set of random amplitudes for each plot presented here: an amplitude $a\in[0,1]$ is sampled afresh for each run. We did not reuse amplitudes in any two executions, nor during the development of the algorithm and tuning of the hyperparameters. As such, BAE is in principle well-fitted to all possible amplitude estimation problems, given its robust performance across hundreds of amplitudes picked at random; and all the demonstrations shown here hold for arbitrary amplitudes. Accordingly, we always assume the absence of prior knowledge about $a$ for BAE, putting it on equal footing with the other algorithms.
Note that the cost of an algorithm is calculated as the total number of queries used throughout its execution. These queries may be split among multiple circuits. This comprehensive definition of cost allows for a direct comparison with the limits of quantum metrology (section 2.5), which is not concerned with the specifics of the resource allocation.
This ends the discussion on the quantities to be plotted. We now discuss how to plot them. Firstly, we will use a double logarithmic scale. We represent the limits in equations (16) and (17) for reference; they take the form of straight lines with 0.5-0.5 and 1-1 slopes respectively, facilitating visual assessments.
Secondly, we customize the y intercept to the datasets to facilitate a visual analysis. Details are provided in Appendix G. Note that the vertical scale offset should still be heeded: it is a relevant cost metric in practice, depending on the precision regime. It may be undesirable to have algorithms requiring a higher quantum resource count from the outset, even if the scaling is Heisenberg-limited. Eventually, the performance of such an algorithm will beat that of another with a smaller offset but a less favorable evolution pace, but whether this happens depends on the target resolution.
Another challenge arises from the use of adaptivity. In quantum metrology, the performance metrics tend to respect average results, due to statistical noise. These averages are straightforward in the case of deterministic algorithms, but not with adaptive ones. In that case, due to a mismatch in x-coordinates, brute force averages would require discarding most data points and using a disproportional amount of executions, growing exponentially with the maximum query number. To avoid this, we find and employ a good approximation that allows us to use all data points while still depicting the statistic reliably. We analyzed multiple possible strategies, and found the best performing one to be binning the points according to their x-coordinates and averaging both their coordinates independently.
Results for mock data are displayed in Figure 3. These data were generated to reproduce ideal Heisenberg-limited behavior, so that the results can be verified. By means of the aforementioned strategy, the same-colored points contained in each bin are summarized by a single point: the star markers. Ideal processing would have these markers lying unbiasedly on the dashed line. Details can be found in Appendix E; Appendix F shows the results using other strategies, demonstrating that our results hold for any processing strategy that behaved well for dummy data.
Figure 3: Behavior of the adopted strategy to process adaptive algorithm data: binning and averaging. Here it is applied to dummy data, which is grouped into bins (different colors) and then summarized (star markers). It comes close to ideal behavior - markers lying on the line.
4.2 Quantum simulation
We want to numerically simulate BAE, other hybrid approaches to QAE, and the canonical QAE algorithm (to be used as a reference). For testing purposes, the quantum parts of each of these algorithms can be simulated efficiently using analytical calculations and multinomial sampling. This does not mean that the circuits can be efficiently simulated by a classical computer, as generating these data requires knowing the value of $a$. Refer to Appendix H for more details.
In addition to the ideal behavior, we want to observe the behavior of the algorithms under the influence of extrinsic noise (i.e. not shot noise). We thus augment equation 9 with an extra parameter $T$ - a coherence time -, and assume an exponential decay affecting the basis states equally. The factor that is independent of $\theta$ assures this symmetry (and proper normalization). The result is expression 4.2.
$P(\psi_{1}\mid m)=e^{-\nicefrac{{m}}{{T}}}\sin^{2}((2m+1)\theta)$
$+\frac{1-e^{-\nicefrac{{m}}{{T}}}}{2}.$ (19)
For practicality, we define a time unit as the time taken by one application of the Grover operator. $T$ can then be expressed using these units. Having a closed expression given by equation 4.2, sampling noise can then be introduced as usual.
5 Results
This section graphically presents the results of testing our BAE algorithm as described in section 4. The statistics are taken over 100 executions. In the comparative plots, we consider the evolution of the median errors to facilitate the visual analysis. Average results are provided in Appendix F for completeness.
Shot noise is present in all tests; in subsection 5.1 no other sources of noise are considered, whereas in section 5.2 finite coherence times are imposed. Section 5.3 presents a comparative analysis, benchmarking BAE against the state-of-the-art.
The code and datasets used for these simulations are publicly available on GitHub [42], along with a demonstrative notebook.
5.1 Noiseless performance of BAE
Figure 4(a) shows the numerical simulation results for BAE. We can see that the evolution of the error is smooth and parallel to the Heisenberg limit.
(a) Baseline.
(b) With additional cost-sparing measures.
Figure 4: Performance of Bayesian amplitude estimation in ideal conditions (shot noise only). We observe Heisenberg-limited estimation. We find extra cost-sparing measures that barely affect performance, while decreasing classical processing costs and increasing quantum parallelism.
Figure 4(b) shows the results of employing the cost sparing measures mentioned in Appendix C: the resampling steps of sequential Monte Carlo were suppressed for utility calculations, and $10$ shots were realized for each $m$. The offset is only slightly worse, and we still observe Heisenberg-limited estimation. The runtime for the classical processing was cut by a factor of 5. In all other simulations, we employ these cost sparing measures.
Note that a slightly better cost offset may be obtained in exchange for less regular behavior, by reducing the number of warm-up shots and shots per measurement, or relaxing the criteria for triggering window expansions (appendices C and J).
We also test annealed BAE, as per subsection 3.1. Figure 5 shows the results, with the estimation still parallel to the Heisenberg limit. The cost offset is similar to BAE under the same circumstances and for unidimensional estimation. This is promising, given that its advantages are expected to show for higher dimensional problems. For instance, multi-parameter noise models may give rise to multimodality, in which case this approach is expected to outpeform the original BAE, which is variance-based.
Figure 5: Performance of annealed Bayesian amplitude estimation in ideal conditions (shot noise only). This variant, despite optimizing statistical representativity rather than uncertainty, yields similar results to the original, and can replace it when variance-based methods are problematic.
5.2 Performance of BAE with decoherence
To demonstrate BAE’s potential for estimation in the presence of noise, we run the same simulations considering a depolarizing channel with a finite coherence time. For these tests, we assume random coherence times $T_{c}\in[2000,5000[$, with the time units described in section 4.2. The time is picked at random for each run and assumed constant within it. BAE uses $500$ shots to estimate the coherence time.
Figure 6 shows the estimation process. We can see that the evolution is initially parallel to the Heisenberg limit, and slows down as the number of queries increases, achieving an intermediate slope between the quantum and classical limits.
Figure 6: Performance of Bayesian amplitude estimation in the presence of decoherence. The estimation is initially Heisenberg limited and then slows down, continuing at a pace between the quantum and classical limits.
In the noiseless case, as the error shrinks, deeper circuits are chosen to improve the estimation error optimally. In this case, where the coherence is limited, there is a trade-off: deeper circuits are more affected by decoherence. Thus, BAE chooses the circuit depth that achieves the best balance between these two considerations, retaining part of the quantum advantage despite the noise.
5.3 Comparative analysis
Finally, we test other algorithms proposed in the literature and compare them with BAE. The results are presented in Figure 7.
(a) Ideal case (shot noise only). BAE achieves the lowest cost for any error, and the lowest error for any cost.
(b) Noisy case (decoherence). BAE achieves the lowest error, as the other algorithms stagnate due to noise.
Figure 7: Performance of quantum amplitude estimation algorithms. Bayesian amplitude estimation (BAE) outperforms all others in both cases. Acronyms are described in Section 2.6; methodology and units are discussed in Section 4.
Starting by the ideal case (Figure 7(a)), we observe that BAE has competitive performance in terms of both complexity and offset: its learning rate (slope) matches that of the best-performing methods, while its offset is better than any other. BAE not only saturates the Heisenberg limit, it also manages to shrink the error faster than any other algorithm as a function of the resource count. This illustrates how even though it can adapt to noisy devices, BAE is not exclusively a NISQ algorithm: it is capable of achieving full quantum advantage in a fault-tolerant scenario.
Turning to the noisy case (finite coherence - Figure 7(b)), BAE remains the best performing algorithm [^1] QAES is not represented because it doesn’t terminate (as discussed in section 2.4). The original algorithm for QAE is not represented due to the runtime, but it is easy to see that it wouldn’t fare well. Reaching $N_{q}=10^{6}$ with the phase estimation based circuit would require a Fourier order $m=19$, and thus $2^{18}$ consecutive applications of the (controlled) Grover operator followed by the QFT. This far exceeds the coherence times., with an even bigger separation.
Among all tested algorithms, BAE achieves the lowest estimation error for any given number of queries. In fact, it achieves the lowest estimation error regardless of cost, as the other algorithms do not seem capable of achieving error below $10^{5}$ due to stagnation in the learning rate. BAE performs reliably even for higher numbers of queries, corresponding to longer executions and higher precision, whereas others with similar costs come to a stall or erratic behavior due to the noise.
Of the other algorithms, some only reveal this trend for higher numbers of queries: FAE, SAE and MLAE (LIS). This is due to their larger offsets and/or slower learning rates; they use shorter circuits than would be optimal, which inflates the quantum cost and delays the effect of decoherence. This does not constitute an advantage: we are interested in the effect of noise for a given magnitude of the error, not of the cost. (Otherwise, one might ”improve” an algorithm’s noise resilience by inflating the cost for no benefit, causing a shift to the right in the plot.) For similar reasons, MLAE-LIS runs into runtime and numerical stability issues for high query numbers, due to the optimization of products envolving large numbers of small terms (likelihoods for tens of thousands of circuits).
Relatedly, we note that the total number of queries may not be straightforwardly related to the impact of decoherence, nor is it the sole predictor of unstable behavior in the presence of noise. These queries are cumulative, and not necessarily proportional to the maximum circuit depth (although they are approximately so for some algorithms, such as the geometric progression circuits of MLAE-EIS).
These are the main results of our work. We have demonstrated that BAE:
  1. 1.
    Achieves Heisenberg-limited estimation in the absence of noise, with a slope as good as the best algorithms;
  2. 2.
    Has a reduced offset as compared to all other tested algorithms;
  3. 3.
    Can, unlike the other algorithms, adapt to noisy scenarios.
The last point is related to other perks coming from the pliable and wide-ranging framework: BAE is capable of accomodating experimental limitations or preferences, assessing the merit of noise models, leveraging previously available information, and exploiting various trade-offs.
6 Conclusions and future work
We propose BAE, a QAE algorithm that is capable of attaining Heisenberg-limited estimation; is highly customizable and capable of negotiating trade-offs across the multiple costs involved; is parallelizable and scalable; and is resilient to noise. Numerical simulations show that our algorithm has a robust performance as compared to state-of-the-art algorithms, both in the presence and in the absence of noise. This makes it especially interesting in the transition between the NISQ and fault tolerant eras, as an algorithm that can interpolate between those regimes.
In particular, BAE is capable of characterizing noise and self-adapting accordingly. This makes it better suited to faulty quantum devices than algorithms that fail to account for extrinsic noise sources. While noise can still slow down the learning rate, properly treating it can minimize this slowdown while safeguarding correctness. We observe that our algorithm continues to learn even after others saturate due to noise.
We additionally propose aBAE, an annealed version of BAE which guides the inference based exclusively on a measure of statistical degeneracy. This makes the cost of evaluating the utility function independent of the number of parameters, and helps assure a stable numerical representation in multi-dimensional scenarios, as those brought about by detailed descriptions of noise.
Relatedly, an interesting direction for future work is to test the algorithm using more complex noise models. Such models are expected to benefit the most from the robustness of the presented methods, namely the efficient multi-dimensional sampling and the annealed variant of BAE. This improvement is a key step towards another target for future work: performing quantum amplitude estimation on real quantum devices, rather than relying on numerical simulations.
Other possible research directions include modifications to specific parts of the BAE algorithm, namely the utility function, the control optimization routine, and the numerical representation. This may further improve the performance or reduce classical costs.
Finally, the BAE algorithm could be directly applied to the characterization tasks mentioned in the introduction, with numerous potential applications in quantum technologies - namely superconducting and photonic quantum computing, and sensing.
The authors thank Ernesto Galvão, Antonio Molero and Bruna Salgado for helpful comments. This work is funded by national funds through FCT – Fundação para a Ciência e a Tecnologia, I.P., under the support UID/50014/2023 (https://doi.org/10.54499/UID/50014/2023). A.R. acknowledges support from FCT under PhD grant 2022.12332.BD.
Appendix A Bayesian inference
@@
Figure 2(a) summarizes the inference process applied to QAE. We use vector notation to emphasize sequences. This does not provide a way of choosing $\vec{m}$, which is the main problem in iterative QAE.
Our scheme relies on substancial enhancements as compared to Figure 2(a), with the goal of choosing good values of $m$ while keeping processing costs down. Figure 2(b) presents a high level view of our scheme.
@@
One alternative is to find the mean, which requires integrating over the distribution. Integration methods are discussed in Appendix B. But before, we note that this extra step arises due to one of Bayesian inference’s strongest features. The mean is just a particular case of an expected value we can compute based on data. More generally, we can calculate the expectation for any function of the parameter(s),
@@
In Figure 2(a), it would take the form of an a priori optimization of $\vec{m}$, the sequence of controls, considering expectations over the prior. However, the quality of the predictions is heavily dependent on the accuracy of the distribution; that is, how close it is to reality (a Dirac delta function on the real value). Thus, this choice may be unsatisfactory, especially in the case of uninformative priors.
Instead, we can unfold the inference process into several steps, considering intermediary distributions $\mathbf{P}(\theta\mid D_{0})$, $\mathbf{P}(\theta\mid D_{0},D_{1})$, and so on. The freedom to do so stems from the structure of Bayes’ rule: one may do a bulk update based on a complete dataset, or consider the data incrementally. In the second case, the posterior of one iteration becomes the prior for the next, allowing empirical knowledge to inform the latter’s controls. The inference process can then occur online and adaptively, exploiting all available information at each point. There are additional benefits to this approach (Appendix C).
We adopt this adaptive approach for QAE, as reflected in Figure 2(b). We consider a single datum per iteration, but the process can be done with arbitrary granularity, dividing the measurements into blocks of several and considering a chunk of data per step. There are multiple trade-offs involved. Such details on the experimental design are discussed in Appendix C, where we also introduce problem-tailored tricks to cut back on the optimization cost.
@@
This finalizes the main outline of the algorithm whose pseudo-code is provided in algorithm 1. The estimator is a construction mapping the posterior to an estimate - such as the mean, the mode or the median.
Appendix B Integration Methods
In section 3 we discussed the strengths of Bayesian inference, and how they are tied to the probabilistic depiction of information. It remains to discuss the specifics of this depiction, which are determinant for the realization of such strengths.